OpenStudy (jkbo):

Please Help Find 2D+A. [1 -1] and D= [1 0] A= [3 -2] [0 1]

OpenStudy (science0229):

First, evaluate 2D\[2D=2*\left[\begin{matrix}1 & 0 \\ 0 & 1\end{matrix}\right]=\left[\begin{matrix}2 & 0 \\ 0 & 2\end{matrix}\right]\]

OpenStudy (science0229):

Then add A to it\[A+2D=\left[\begin{matrix}1 & -1 \\ 3 & -2\end{matrix}\right]+\left[\begin{matrix}2 & 0 \\ 0 & 2\end{matrix}\right]=\left[\begin{matrix}1+2 & -1+0 \\ 3+0 & -2+2\end{matrix}\right]=\left[\begin{matrix}3 & -1 \\ 3 & 0\end{matrix}\right]\]

OpenStudy (jkbo):

thank you very much

OpenStudy (science0229):

Welcome :)